# Learning Regularization for Graph Inverse Problems

### v1

@@ 5 Experiments
We evaluate the five models discussed in Section 4 on the diverse set of GRIPs discussed in Section 2. We utilize several datasets to demonstrate GRIPs, ranging from synthetic graphs (CLUSTER) to geometric datasets (ShapeNet), natural images (CIFAR10), road traffic (METR-LA), and spread of disease (Chickenpox-Hungary, shortened to CPOX). We provide full details on the datasets and our motivation for using them to demonstrate different problems, in Appendix D. We also provide a comprehensive description of the training, evaluation, and hyperparameter selection procedures for each problem in Appendix E, followed by a discussion of the complexity and runtimes of the methods, in Appendix G. Our code is implemented in Pytorch (Paszke et al. 2017).
@@ 5.1 Property Completion
In the Property Completion problem, defined in Equation (2) and illustrated in Figure 1, for each graph, we randomly sample $nb$ nodes per class, and the selected nodes’ labels are the observed data. For example, in Figure 5, we show an example where the property to be recovered are the node labels and only $nb$ per class of them are available as input, besides possible meta-data such as coordinates. We present results with $nb=4$ in Table 1, and results on $nb=8,16$ in Appendix F. We observe an accuracy increasing trend is evident when increasing $nb$, and a consistently better performance offered by neural approaches.
Also, we study the effect of meta-data, as shown in Table 2 (and in Appendix F), where a significant drop in performance is seen in the models when meta-data was not used.
@@ 5.2 Inverse Source Estimation
We now discuss the results for the Inverse Source Estimation problem, which is described by the forward operator in Equation (3) and Figure 2. In this problem, node states x are diffused by $k$ applications of the adjacency matrix where $k=4,8,16$, and their diffused result is the observed data. As $k$ grows, the problem becomes harder. The results with $k=16$ are reported in on the CLUSTER dataset in Table 3, and on the CPOX dataset in Table 4, with additional results in Appendix F. Our results indicate that all neural methods achieve better performance than classical methods.
@@ 5.4 Edge Property Recovery
We study a nonlinear inverse problem, compared with the previously presented linear problems. Here, the goal is to recover edge weights given data that was diffused using these weights, as described in Equation (5) and by Figure 4. The results are given on the superpixels CIFAR10 dataset, in Table 6, where we see that classical methods struggle to offer a low relative error, compared with 55~{}55 times lower (better) error given by Var-GNN and ISS-GNN. The results on different GRIPs and datasets show the generality of the framework.
@@ A.1 Solving Inverse Problems with Neural Networks
Inverse problems are a crucial area in many scientific and engineering domains, where one aims to infer the underlying causes from observed effects. Traditional methods for solving inverse problems include variational approaches and regularization techniques, such as Tikhonov regularization, Bayesian inference (Tikhonov and Arsenin 1977; Kaipio and Somersalo 2006) and inverse scale space techniques (Burger et al. 2006). These methods often rely on explicit mathematical formulations and assumptions about the underlying models. In recent years, neural networks have emerged as a powerful tool for addressing inverse problems, owing to their ability to learn complex mappings from data. Starting from the work (Adler and Öktem 2017), which used proximal gradient descent, to related work (Mardani et al. 2018; Aggarwal, Mani, and Jacob 2018; Bai et al. 2020; Ongie et al. 2020; Jin et al. 2017), which used similar algorithms, deep networks have been used as priors and leveraged the powerful feature extraction capabilities to enhance reconstruction accuracy. Recent work used a variational framework (Eliasof, Haber, and Treister 2023) to achieve similar goals. More recently, diffusion models have been used to construct a prior for inverse problems (Chung et al. 2022). While at inference, they perform very similar procedures as in the methods proposed above, they are trained very differently and not in an end-to-end fashion. We refer the reader for a comprehensive review of neural approaches for solving inverse problems to Arridge et al. (2019)
@@ A.2 Inverse Problems on Graphs
Compared to the solution of inverse problems on structured grids (e.g., 2D deblurring or 3D tomographic reconstruction), there is a much smaller body of work for the solution of inverse problems on graphs, which are unstructured in nature. Traditional methods for graph inverse problems, do not include learnable networks and typically rely on optimization techniques and probabilistic models. Similar to standard inverse problems, one of the foundational approaches in this domain is the use of graph regularization techniques, which incorporate prior knowledge about the graph structure to solve inverse problems more effectively (Zhou et al. 2004; Zhou and Schölkopf 2005). These methods have been widely used in semi-supervised learning and signal processing on graphs. In recent years the application of GNNs to inverse problems have been tailored to a particular application. Two recent cases are node classification (see e.g. (Veličković et al. 2018) and reference within) and source estimation (Yan, Fang, and He 2023; Huang et al. 2023). In recent work the problem of source estimation have been solved by using diffusion models as priors (Ling et al. 2024). This idea is similar to the use of diffusion models as priors in the context of inverse problems as discussed above.
@@ Appendix B Broader Impact and Ethical Statement
In this paper we proposed a framework for solving GRIPs, and we demonstrated it on several key GRIP instances, on several difference datasets. Combining understanding for learned regularization to solving inverse problems, together with GNNs, can have a positive potential impact on the research of GNNs as well as their applicability to solve various real-world problems. We are not aware of negative societal impacts that can arise from the research presented in this paper.
@@ Appendix C Architecture Details
In order to obtain a fair evaluation of the methods, as use the same GNN backbone, such that the only difference between Var-GNN, ISS-GNN, and Prox-GNN is their overall iteration definition, not the learned regularizer backbone. Specifically, we employ a simple GCN (Kipf and Welling 2016) with a specified number of layers, as later discussed in Appendix E.2. In the case of edge recovery problem, since the property to be recovered is an edge property rather than a node one (as in all the other problems), we use a similar approach as is done with the Link Prediction task (Bronstein et al. 2021), and concatenate the final node features according to the given connectivity, and employ an additional MLP to predict the edge property. While other GNN backbones can be studied for specific problems, such as GAT (Veličković et al. 2018) or GIN (Xu et al. 2019), this is not the main contribution of this paper. Instead, it is on proposing a general framework for solving GRIPs with GNNs.
@@ Appendix D Datasets
CLUSTER. Is a dataset introduced in (Dwivedi et al. 2022), and it consists of synthetic graphs where nodes are grouped into distinct clusters, and the objective is to accurately predict the cluster membership of each node. This is classification problem. The CLUSTER dataset includes a total of 6 communities and therefore 6 classes for nodes. The graphs are of sizes $40$-$190$ nodes. The target labels for nodes are defined as the community/class labels. This dataset was used for the GRIP of graph property completion. During model training, a fixed node budget per class label per graph $nb$ is set, where $nb$ random nodes per label per graph were made visible for training with other nodes masked. During evaluation on the test set, the same node budget $nb$ is kept for the test set. The graph property completion problem is suited for this dataset as it seeks to solve the GRIP of recovering class labels of nodes during inference on the test set, which aligns with the original intent for this dataset of node classification. The dataset contains 10K training samples and 1K test samples. This dataset contains no additional meta-data ,i.e., no node features.
@@
ShapeNet. The ShapeNet part level segmentation (Yi et al. 2016) dataset offered by PyTorch Geometric, is a subset of the larger ShapeNet dataset. It includes point clouds in 16 categories with 50 classes, with each category representing a distinct type of 3D object with the class of a point being a part annotation. The dataset is commonly used for 3D object classification and part segmentation tasks. There are $16,881$ objects in total represented by $2048$ points each, from which we sample $1024$ points uniformly. In all experiments, graphs were generated with the K-Nearest-Neighbours algorithm with $k=10$, and using the same splits as in Wang et al. (2018). The associated meta-data are the normal vectors, and the 3D ($xyz$) coordinates for respective points. The associated node states ${\bf x}$ are their class labels. This dataset was chosen for the graph completion problem due to the availability of meta-data which could be utilized to aid in solving the problem, and to study the effect on performance when meta-data was excluded.
@@
METR-LA. The METR-LA dataset (Jagadish et al. 2014) consists of traffic data on a road network graph, where nodes represent traffic sensors, and edges represent the road network connectivity between the sensors. The dataset comprises 34,272 samples of the road network traffic speed at different times, at 207 nodes (sensors). The available meta-data includes time of measurement. We follow the data splits presented in Taheri and Berger-Wolf (2019). This dataset is commonly used in regression problems where the objective is to predict traffic at future time points. In our experiments however, we do not perform prediction tasks and use the static graphs as-is for the inverse graph transport problem. For an example problem, consider the case where the average speed of vehicles in a road network was known for various segments, and a problem of interest was to infer the speeds of vehicles at specific points. The average speed is the sum (divided by the number of segments) of the speeds measured by sensors. This dataset fits into the framework of such a problem.
@@
Chickenpox Cases in Hungary (CPOX). This dataset was introduced in Rozemberczki et al. (2021), and it consists of weekly chickenpox cases recorded in Hungary, represented on a graph where nodes are counties, and edge weights represent the spatial proximity between the counties. There are 520 time samples, measured weekly. Usually, CPOX is used for temporal prediction tasks, where node features are lagged weekly counts of the chickenpox cases. Here we use it differently by setting lag to 00, effectively making the target node states, ${\bf x}$, the number of chickenpox cases recorded in each county at the current week. This dataset was used for the inverse graph transport problem, which is a regression task. The available meta-data is the week of measurement of chickenpox cases, and the edge weights are the distances between counties. We follow the splitting approach in Rozemberczki et al. (2021). The motivation behind the use of this dataset for the task of inverse source estimation is as follows: it is conceivable that it is of interest to policy makers to know the infection spread patterns across various towns (nodes), and that the available data is only the observed infection counts in towns after the spread has commenced. The GRIP would then be to infer the initial infection counts ${\bf x}$ of towns in order to pinpoint the source of the epidemic.
@@
CIFAR10 (superpixels). This dataset was introduced in (Dwivedi et al. 2022), and it includes superpixel images of the CIFAR10 dataset. Each graph includes positions of the superpixels, their RGB value, as well as the class label of the image. To define edge weights to be recovered in the edge property recovery task, a regression task, we define a random connectivity to the nodes of the superpixels using the Erdos-Renyi distribution with edge probability of 5% for each graph in a pre-processing step. The graph remains identical throughout the epochs and evaluation, such that results are deterministic. Then, we compute the edge weight to be recovered as the inverse $\ell_{2}$ distance of the connected nodes, according to their position in the image. The length of the history vector ${\bf d}^{\rm obs}$ in our experiments (results provided in Table 6) was set to $1$ the most difficult setting - more historical samples ($>1$) would add more information to make the problem easier.Different to the CLUSTER dataset, there is a geometrical interpretation of the superpixels with their positions and edge weights (inverse $\ell_{2}$ distance of the connected nodes), and so this GRIP seeks to recover the distance between the superpixels of an image.
@@ Appendix E Experimental Settings
The experiments were run on an Nvidia RTX A6000 GPU with $48$ GB of memory. Each experiment was run 3 times with seeds $0,1,2$ to compute the mean and standard deviation of the performance of the different models. For GRIP experiments involving the ShapeNet and CLUSTER datasets the evaluation metric was Accuracy (%), with the data fits being assessed using the Cross-Entropy (CE) loss between the observed data ${\bf d}^{\rm{obs}}$ (one-hot encoded class labels) and the reconstructed data $F(F^{{\dagger}}({\bf d}^{\rm obs}))$ (the logits predicted by the models). For GRIP experiments involving the METR-LA and CPOX datasets, the evaluation metric was the normalized mean squared error (nMSE). We elaborate more on these for clarity in E.1.
@@ E.2 Hyper-parameter Details
For each dataset, task, and setting, the optimal hyper-parameters were found using Bayesian search on the most difficult setting for the given task. These optimal hyper-parameters were then used to train and test the other settings for the model for the task being considered. For example a hyper-parameter search sweep was conducted for $nb=4$ for the CLUSTER graph completion task in Table 1 for ISS-GNN and the optimal hyper-parameters were used to train ISS-GNN models for $nb=8,16$. Similar sweeps were conducted for the other models. The exceptions to this were the CPOX dataset in Table 17 for which a hyper-parameter sweep was carried out for Prox-GNN for each diffusion step setting ($k=4,8,16$), and the CIFAR10 dataset for which we used the hyper-parameters from the ShapeNet graph completion experiments.
For reference, the hyper-parameter search space for the various hyper-parameters are given in Tables 7, 8 and 9. The search space for the batch size during training was {8, 16, 32, 64} for all datasets.
@@ Appendix F Additional Experimental Results
Tables 15, 17, 16 and 18 show the full range of our numerical experiments, including additional results with different settings of the GRIPs.
In Table 15, ISS-GNN is seen to exhibit decreasing performance with increasing diffusion steps $k$, whilst Var-GNN seems to exhibit consistent performance across $k$ values. At all diffusion steps however, ISS-GNN and Var-GNN exhibit significant performance over Prox-GNN. All neural approaches show significant improvements from the baseline performances.

### v2

@@ 5 Experiments
We evaluate the five models discussed in Section 4 on the diverse set of GRIPs discussed in Section 2. We utilize several datasets to demonstrate GRIPs, ranging from synthetic graphs (CLUSTER) to geometric datasets (ShapeNet), natural images (CIFAR10), road traffic (METR-LA), and spread of disease (Chickenpox-Hungary, shortened to CPOX). We provide full details on the datasets and our motivation for using them to demonstrate different problems, in Appendix D. We also provide a comprehensive description of the training, evaluation, and hyperparameter selection procedures for each problem in Appendix E, hyperparameter sensitivity results in F, and a discussion of the complexity and runtimes of the methods, in Appendix G. Our code is implemented in Pytorch (Paszke et al. 2017), and is available at https://github.com/nafi007/GraphInverseProblems.
@@ 5.1 Property Completion
In the Property Completion problem, defined in Equation (2) and illustrated in Figure 1, for each graph, we randomly sample $nb$ nodes per class, and the selected nodes’ labels are the observed data. For example, in Figure 5, we show an example where the property to be recovered are the node labels, and only $nb$ per class of them are available as input, besides possible meta-data such as coordinates. We present results with $nb=4$ in Table 1, and results on $nb=8,16$ in Appendix F. We observe an accuracy-increasing trend is evident when increasing $nb$ and a consistently better performance offered by neural approaches. Also, we study the effect of meta-data, as shown in Table 2 (and in Appendix F), where a significant drop in performance is seen in the models when meta-data was not used.
@@ 5.2 Inverse Source Estimation
We now discuss the results for the Inverse Source Estimation problem, which is described by the forward operator in Equation (3) and Figure 2. In this problem, node states x are diffused by $k$ applications of the adjacency matrix where $k=4,8,16$, and their diffused result is the observed data. As $k$ grows, the problem becomes harder. The results with $k=16$ are reported in on the CLUSTER dataset in Table 3, and on the CPOX dataset in Table 4, with additional results in Appendix F. We note that in the case of inverse source estimation, our ISS-GNN becomes similar to Huang et al. (2023). Our results indicate that all neural methods achieve better performance than classical methods.
@@ 5.4 Edge Property Recovery
We study a nonlinear inverse problem, differently than the previously presented linear problems. Here, the goal is to recover edge weights given data that was diffused using these weights, as described in Equation (5) and by Figure 4. Results on the superpixels CIFAR10 dataset are shown in Table 6, where we see that classical methods struggle to offer a low relative error, compared with 55~{}55 times lower (better) error given by Var-GNN and ISS-GNN. The results on different GRIPs and datasets show the generality of the framework.
@@ 6 Conclusions and Discussion
Acknowledgments
ME is funded by the Blavatnik-Cambridge fellowship, the Accelerate Programme for Scientific Discovery, and the Maths4DL EPSRC Programme.
@@ A.1 Solving Inverse Problems with Neural Networks
Inverse problems are a crucial area in many scientific and engineering domains, where one aims to infer the underlying causes from observed effects. Traditional methods for solving inverse problems include variational approaches and regularization techniques, such as Tikhonov regularization, Bayesian inference (Tikhonov and Arsenin 1977; Kaipio and Somersalo 2006) and inverse scale-space techniques (Burger et al. 2006). These methods often rely on explicit mathematical formulations and assumptions about the underlying models. In recent years, neural networks have emerged as a powerful tool for addressing inverse problems, owing to their ability to learn complex mappings from data. Starting from the work (Adler and Öktem 2017), which used proximal gradient descent, to related work (Mardani et al. 2018; Aggarwal, Mani, and Jacob 2018; Bai et al. 2020; Ongie et al. 2020; Jin et al. 2017), which used similar algorithms, deep networks have been used as priors and leveraged the powerful feature extraction capabilities to enhance reconstruction accuracy. Recent work used a variational framework (Eliasof, Haber, and Treister 2023) to achieve similar goals. More recently, diffusion models have been used to construct a prior for inverse problems (Chung et al. 2022). While at inference, they perform very similar procedures as in the methods proposed above, they are trained very differently and not in an end-to-end fashion. We refer the reader for a comprehensive review of neural approaches for solving inverse problems to Arridge et al. (2019)
@@ A.2 Inverse Problems on Graphs
Compared to the solution of inverse problems on structured grids (e.g., 2D deblurring or 3D tomographic reconstruction), there is a much smaller body of work for the solution of inverse problems on graphs, which are unstructured in nature. Traditional methods for graph inverse problems, do not include learnable networks and typically rely on optimization techniques and probabilistic models. Similar to standard inverse problems, one of the foundational approaches in this domain is the use of graph regularization techniques, which incorporate prior knowledge about the graph structure to solve inverse problems more effectively (Zhou et al. 2004; Zhou and Schölkopf 2005). These methods have been widely used in semi-supervised learning and signal processing on graphs. In recent years, the application of GNNs to inverse problems has been tailored to a particular application. Two recent cases are node classification (see e.g. (Veličković et al. 2018) and reference within) and source estimation (Yan, Fang, and He 2023; Huang et al. 2023). In recent work, the problem of source estimation has been solved by using diffusion models as priors (Ling et al. 2024). This idea is similar to the use of diffusion models as priors in the context of inverse problems as discussed above.
@@ Appendix B Broader Impact and Ethical Statement
In this paper, we proposed a framework for solving GRIPs, and we demonstrated it on several key GRIP instances on several different datasets. Combining an understanding of learned regularization to solving inverse problems, together with GNNs, can have a positive potential impact on the research of GNNs as well as their applicability to solving various real-world problems. Future directions include the development of theoretical properties of GRIPs and their solutions with GNNs, in particular convergence properties and out-of-distribution behavior. We are not aware of negative societal impacts that can arise from the research presented in this paper.
@@ Appendix C Architecture Details
In order to obtain a fair evaluation of the methods, we use the same GNN backbone, such that the only difference between Var-GNN, ISS-GNN, and Prox-GNN is their overall iteration definition, not the learned regularizer backbone. Specifically, we employ a simple GCN (Kipf and Welling 2016) with a specified number of layers, as later discussed in Appendix E.2. In the case of edge recovery problem, since the property to be recovered is an edge property rather than a node one (as in all the other problems), we use a similar approach as is done with the Link Prediction task (Bronstein et al. 2021), and concatenate the final node features according to the given connectivity, and employ an additional MLP to predict the edge property. While other GNN backbones can be studied for specific problems, such as GAT (Veličković et al. 2018) or GIN (Xu et al. 2019), this is not the main contribution of this paper. Instead, it is on proposing a general framework for solving GRIPs with GNNs.
@@ Appendix D Datasets
CLUSTER. Is a dataset introduced in (Dwivedi et al. 2022), and it consists of synthetic graphs where nodes are grouped into distinct clusters, and the objective is to accurately predict the cluster membership of each node. This is a classification problem. The CLUSTER dataset includes a total of 6 communities and, therefore, 6 classes for nodes. The graphs are of sizes $40$-$190$ nodes. The target labels for nodes are defined as the community/class labels. This dataset was used for the GRIP of graph property completion. During model training, a fixed node budget per class label per graph $nb$ is set, where $nb$ random nodes per label per graph are made visible for training with other nodes masked. During the evaluation of the test set, the same node budget $nb$ is kept for the test set. The graph property completion problem is suited for this dataset as it seeks to solve the GRIP of recovering class labels of nodes during inference on the test set, which aligns with the original intent for this dataset of node classification. The dataset contains 10K training samples and 1K test samples. This dataset contains no additional meta-data,i.e., no node features.
@@
ShapeNet. The ShapeNet part level segmentation (Yi et al. 2016) dataset offered by PyTorch Geometric, is a subset of the larger ShapeNet dataset. It includes point clouds in 16 categories with 50 classes, with each category representing a distinct type of 3D object with the class of a point being a part annotation. The dataset is commonly used for 3D object classification and part segmentation tasks. There are $16,881$ objects in total represented by $2048$ points each, from which we sample $1024$ points uniformly. In all experiments, graphs were generated with the K-Nearest-Neighbors algorithm with $k=10$, and using the same splits as in Wang et al. (2018). The associated meta-data are the normal vectors, and the 3D ($xyz$) coordinates for respective points. The associated node states ${\bf x}$ are their class labels. This dataset was chosen for the graph completion problem due to the availability of meta-data, which could be utilized to aid in solving the problem and to study the effect on performance when meta-data was excluded.
@@
METR-LA. The METR-LA dataset (Jagadish et al. 2014) consists of traffic data on a road network graph, where nodes represent traffic sensors, and edges represent the road network connectivity between the sensors. The dataset comprises 34,272 samples of the road network traffic speed at different times, at 207 nodes (sensors). The available meta-data includes the time of measurement. We follow the data splits presented in Taheri and Berger-Wolf (2019). This dataset is commonly used in regression problems where the objective is to predict traffic at future time points. In our experiments, however, we do not perform prediction tasks and use the static graphs as-is for the inverse graph transport problem. For an example problem, consider the case where the average speed of vehicles in a road network was known for various segments, and a problem of interest was to infer the speeds of vehicles at specific points. The average speed is the sum (divided by the number of segments) of the speeds measured by sensors. This dataset fits into the framework of such a problem.
@@
Chickenpox Cases in Hungary (CPOX). This dataset was introduced in Rozemberczki et al. (2021), and it consists of weekly chickenpox cases recorded in Hungary, represented on a graph where nodes are counties and edge weights represent the spatial proximity between the counties. There are 520 time samples, measured weekly. Usually, CPOX is used for temporal prediction tasks, where node features are lagged weekly counts of the chickenpox cases. Here we use it differently by setting lag to 00, effectively making the target node states, ${\bf x}$, the number of chickenpox cases recorded in each county at the current week. This dataset was used for the inverse graph transport problem, which is a regression task. The available meta-data is the week of measurement of chickenpox cases, and the edge weights are the distances between counties. We follow the splitting approach in Rozemberczki et al. (2021). The motivation behind the use of this dataset for the task of inverse source estimation is as follows: it is conceivable that it is of interest to policy makers to know the infection spread patterns across various towns (nodes), and that the available data is only the observed infection counts in towns after the spread has commenced. The GRIP would then be to infer the initial infection counts ${\bf x}$ of towns in order to pinpoint the source of the epidemic.
@@
CIFAR10 (superpixels). This dataset was introduced in (Dwivedi et al. 2022), and it includes superpixel images of the CIFAR10 dataset. Each graph includes the positions of the superpixels, their RGB value, as well as the class label of the image. To define edge weights to be recovered in the edge property recovery task, a regression task, we define random connectivity to the nodes of the superpixels using the Erdos-Renyi distribution with an edge probability of 5% for each graph in a pre-processing step. The graph remains identical throughout the epochs and evaluation, such that results are deterministic. Then, we compute the edge weight to be recovered as the inverse $\ell_{2}$ distance of the connected nodes, according to their position in the image. The length of the history vector ${\bf d}^{\rm obs}$ in our experiments (results provided in Table 6) was set to $1$ the most difficult setting - more historical samples ($>1$) would add more information to make the problem easier. Different from the CLUSTER dataset, there is a geometrical interpretation of the superpixels with their positions and edge weights (inverse $\ell_{2}$ distance of the connected nodes), and so this GRIP seeks to recover the distance between the superpixels of an image.
@@ Appendix E Experimental Settings
The experiments were run on an Nvidia RTX A6000 GPU with $48$ GB of memory. Each experiment was run 3 times with seeds $0,1,2$ to compute the mean and standard deviation of the performance of the different models. For GRIP experiments involving the ShapeNet and CLUSTER datasets, the evaluation metric was Accuracy (%), with the data fits being assessed using the Cross-Entropy (CE) loss between the observed data ${\bf d}^{\rm{obs}}$ (one-hot encoded class labels) and the reconstructed data $F(F^{{\dagger}}({\bf d}^{\rm obs}))$ (the logits predicted by the models). For GRIP experiments involving the METR-LA and CPOX datasets, the evaluation metric was the normalized mean squared error (nMSE). We elaborate more on these for clarity in E.1.
@@ E.2 Hyper-parameter Details
For each dataset, task, and setting, the optimal hyper-parameters were found using a Bayesian search on the most difficult setting for the given task. These optimal hyper-parameters were then used to train and test the other settings for the model for the task being considered. For example, a hyper-parameter search sweep was conducted for $nb=4$ for the CLUSTER graph completion task in Table 1 for ISS-GNN, and the optimal hyper-parameters were used to train ISS-GNN models for $nb=8,16$. Similar sweeps were conducted for the other models. The exceptions to this were the CPOX dataset in Table 17 for which a hyper-parameter sweep was carried out for Prox-GNN for each diffusion step setting ($k=4,8,16$), and the CIFAR10 dataset for which we used the hyper-parameters from the ShapeNet graph completion experiments.
For reference, the hyper-parameter search space for the various hyper-parameters is given in Tables 7, 8 and 9. The search space for the batch size during training was {8, 16, 32, 64} for all datasets.
@@ Appendix F Additional Experimental Results
Tables 15, 16, 17 and 18 show the full range of our numerical experiments, including additional results with different settings of the GRIPs.
In Table 15, for the inverse source estimation problem, ISS-GNN is seen to exhibit decreasing performance with increasing diffusion steps $k$, whilst Var-GNN seems to exhibit consistent performance across $k$ values. At all diffusion steps, however, ISS-GNN and Var-GNN exhibit significant performance over Prox-GNN. All neural approaches show significant improvements from the baseline performances.
@@
Lastly, we provide an exemplary hyperparameter sensitivity study in Table 19, starting from the best setup: cglsIter=8, channels=64, layers=16, lr=0.007, solveIter=8, wd=5e-06, and varying one parameter at a time, on the ShapeNet property completion task with nb = 4 for Var-GNN. The results demonstrate that the method is relatively stable with respect to the chosen hyperparameters.